### v1

@@ 4.1 Design
The permutation of indexes changes at every device instantiation, so that the mapping is not static. The size of these in-memory supporting data structures (array and bitfield) is $O\left(\frac{N}{S_{P}}\log\frac{N}{S_{P}}\right)$, as usual. The lazy allocation technique is also what allows us to overcommit the total physical storage space: we can have the sum of the sizes of the logical volumes exceed the total physical storage space, as long as the sum of “actually used” spaces does not. However, it might be useful to intentionally limit this overcommitment (for the opened volumes) to decrease the risk of I/O errors and improve user experience. This can be done using the metadata field in the VMB ciphertexts, as explained in Section 6.5.
Algorithm 3 NewSlice($V_{i},\;\sigma$)
1:while $\mathtt{bfld}[\mathtt{prmslices}[\mathtt{octr}]]=\mathtt{occupied}$ do: $\mathtt{octr}:=\mathtt{octr}+1$
2:$\varPsi_{new}\leftarrow\mathtt{prmslices}[\mathtt{octr}]$
3:$\mathtt{bfld}[\mathtt{prmslices}[\mathtt{octr}]]:=\mathtt{occupied}$
4:$\mathtt{PosMap}_{i}[\sigma]:=\varPsi_{new}$
5:return $\varPsi_{new}$
4.1.3 Volume operations.
@@ 4.2 Operational Model
A simple observation shows how a legitimate-looking usage mode of Shufflecake actually entails a high risk of data corruption. If we do not open all $\ell$ existing volumes, and instead only open the ones we plan to use, we do not load all the slice maps in RAM, which leads to an incorrect reconstruction of the complete device’s bitfield bfld of free physical slices. The physical slices belonging to the still-closed volumes will be counted as free, and will therefore possibly be allocated to the open volumes during data write, which would then overwrite their content. This can only be avoided with certainty by always opening every volume, regardless of which ones we are going to use: if the password to a volume is not provided, Shufflecake has no way of detecting its existence. It then follows from the overcommitment of the physical storage space that we risk re-using its physical slices for some other volume.
However, mitigation does need to be perfect. It could be possible to reduce the risk of corruption when not opening all volumes by using some form of error-correction on the unopened volumes (hence sacrificing some space), and then trying to recover the volume if corruption happens. We discuss this in Section 4.2.1.
@@
In normal operating conditions, when not confronted with the adversary, the recommended course of action for the user is to unlock all the volumes present on the device, in order to prevent data corruption as explained before. The reason behind the design choice of chaining the volumes into a linked list to help the user in that regard: this way, the user is able to open all volumes by just providing the password to $V_{\ell}$. In our implementation, this is actually the mandated semantic of the open operation: the user only provides the password of the last volume they want to open, the previous ones are automatically opened. Other implementations are of course free to ignore the $\mathtt{\mathsf{VMK}}_{i-1}$ field in the volume header, and give more flexibility to the user, if aware of the risks entailed.
@@
When questioned by the adversary and forced to reveal passwords, the user must obviously not surrender more than $\ell-1$ of them (otherwise there would be nothing left to protect). Although irrelevant for the cryptographic security of the scheme, we stress that, in order for the user’s lie to be credible, they must only reveal the decoy passwords under a certain amount of pressure, or after some time has passed. Notice that a responsible and safe use of Shufflecake puts on the user the burden of being able recall quickly and reliably these decoy passwords, even under distress. This might be hard to get right given that the recommended course of action for the user in daily use is to only open the most hidden volume. It is up to the user to define the maximum number of volumes $\ell$ that makes them comfortable in this task. Shufflecake implementations might include additional features to aid the user in this sense, for example a function to check the password of a decoy volume without actually opening it, or even a puzzle which, with some random probability, prompts the user to also insert the password for a decoy volume when opening a hidden one.
@@
As we previously discussed, one big operational difference between Shufflecake (but also other solutions like, e.g., HiVE) and TrueCrypt is that with Shufflecake “the adversary does not know when she can stop questioning you”, because there is no way to prove that a given password unlocks all the existing volumes on a given device (unless it’s the password unlocking the max-th volume). In TrueCrypt, instead, there is either just a regular volume, or a decoy and a hidden volume. This distinction might be important in those scenarios where a user wants to avoid the possibility of looking uncooperative to a certain adversary. In such scenarios, the user might want to have the choice of surrendering all the volumes, and a method to convince the adversary that no other volumes exist, possibly using an additional “full disclosure password” that we will call safeword.
One simple way to implement this, even in Shufflecake, is to actually create all max volumes when initializing a device. Clearly, remembering max passwords would be quite cumbersome for the user, so the solution is to only remember $\ell+1$ passwords instead: those for the $\ell$ volumes that are actually desired, and the extra one for the last max-th volume, which is going to be the safeword. In fact, the user will never need to open more than $\ell$ volumes for regular use. As discussed in Section 4.2, this might harm the consistency of the other, unopened volumes, but this is not important: by using the safeword, the user would still be able to convince the adversary that all volumes have been revealed due to the linkage between them and the ciphertext authentication in the headers. Analogously, in solutions like TrueCrypt, a simple way to implement a safeword is to actually always create a hidden volume, even if only a standard volume is desired.
@@ 4.3 Security
  • The symmetric encryption scheme will be replaced by an ideal cipher $\mathcal{E}$, mapping $\lambda$-bit keys and IVs to truly-random permutations over $\{0,1\}^{\lambda}$.
@@
Let us consider Experiment 2.1 under the constraints explained in Section 2.3.3, and let $D$ be the “challenge disk snapshot” provided to the adversary $\mathcal{A}$ by the game (i.e., either $D_{0}$ or $D_{1}$ according to the secret bit $b$). For the given $\mathcal{A}$, we will consider $\ell$, the decoy passwords $P_{1},...,P_{\ell-1}$, and the access patterns $O_{0}$ and $O_{1}$ as public parameters of the game instance.
@@
Then, let us define $Q$ to be the ordered sequence of all queries $\{q_{i}\}_{i}$ made by $\mathcal{A}$ to the oracles after receiving the challenge $D$, and define $n:=|Q|$. Also define $Q_{i}$ to be the sequence of queries $(q_{1},\ldots,q_{i})$ up to query $q_{i}$. Analogously, let us define $R$ to be the ordered sequence of all responses $\{r_{i}\}_{i}$ returned by the oracles; also define $R_{i}$ to be the sequence of responses $(r_{1},\ldots,r_{i})$ up to response $r_{i}$. We consider the execution of $\mathcal{A}$ as the execution of a sequence of single-query stateful adversaries, where the state is just the ‘history’ of the previous queries: $\mathcal{A}_{1}\left(D\right)\to q_{1}$, $\mathcal{A}_{2}\left(D,Q_{1},R_{1}\right)\to q_{2}$, and so on until $\mathcal{A}_{n-1}\left(D,Q_{n-1},R_{n-1}\right)\to q_{n}$ and $\mathcal{A}_{n}\left(D,Q_{n},R_{n}\right)\to b^{\prime}$.
Let $\mathsf{KEK}_{i}$, $\mathsf{VMK}_{i}$ and $\mathsf{VEK}_{i}$ be, respectively, the key-encryption key, the VMB key, and the volume encryption key of volume $V_{i}$. Rigorously speaking, in the security game, the values $P_{\ell},\mathsf{KEK}_{\ell},\mathsf{VMK}_{\ell}$, and $\mathsf{VEK}_{\ell}$ are only sampled if $b=0$. Let us instead consider them to be sampled anyway, and left unused in the case $b=1$. Let us denote by $\mathcal{S}$ the tuple $(P_{\ell},\mathsf{KEK}_{\ell},\mathsf{VMK_{\ell}},\mathsf{VEK}_{\ell})\in\{0,1\}^{4\lambda}$. Let us define the event $E_{i}$ as the event that either of $P_{\ell}$, $\mathsf{KEK}_{\ell}$, $\mathsf{VMK}_{\ell}$, or $\mathsf{VEK}_{\ell}$ appear in query $q_{i}$ (we say that query $q_{i}$ strikes). Finally, let us define $E:=E_{1}\cup\ldots\cup E_{n}$ the event that at least one query strikes. We will first prove two lemmata.
@@
$\underset{}{\mathsf{Pr}}\left[{E}\right]=\underset{}{\mathsf{Pr}}\left[{E\,|\,b=0}\right]=\underset{}{\mathsf{Pr}}\left[{E\,|\,b=1}\right]=\mathsf{negl}(\lambda)$. (The adversary can only guess one of the secrets of $V_{\ell}$ with negligible probability.)
@@
This tuple gets sampled uniformly at random from $\{0,1\}^{4\lambda}$, so its distribution does not depend on the public values. Since the oracles implement ideal cryptographic primitives, it follows that their outputs are statistically independent from their inputs; this does not just hold marginally for single input-output pairs, but jointly: any tuple of inputs is statistically independent from the tuple of corresponding outputs (for the ideal cipher $\mathcal{E}$, this only holds for the key inputs). In particular, $\mathcal{S}$ is statistically independent from the whole tuple $(D,R)$. Since $\mathcal{A}_{i}$’s inputs, namely the tuple $(D,Q_{i-1},R_{i-1})$, are a (randomised) function of $(D,R)$, we deduce that its output, namely $q_{i}$, must also be independent of $\mathcal{S}$.
Therefore, since $(P_{\ell},\mathsf{KEK}_{\ell},\mathsf{VMK}_{\ell},\mathsf{VEK}_{\ell})$ are uniform, the probability that $q_{i}$ contains, say, $P_{\ell}$, is $2^{-\lambda}$. Thus, by the union bound, $\underset{}{\mathsf{Pr}}\left[{E_{i}}\right]\leq 4\cdot 2^{-\lambda}$. Using the union bound again, we get:
$\underset{}{\mathsf{Pr}}\left[{E}\right]\leq\sum_{i=1}^{n}\underset{}{\mathsf{Pr}}\left[{E_{i}}\right]\leq n\cdot 4\cdot 2^{-\lambda}$
@@
$\underset{}{\mathsf{Pr}}\left[{\mathcal{A}\left(D\right)\!\to\!1|b\!=\!0\wedge\bar{E}}\right]=\underset{}{\mathsf{Pr}}\left[{\mathcal{A}\left(D\right)\!\to\!1|b\!=\!1\wedge\bar{E}}\right]$. (Unless she can guess one of $V_{\ell}$’s secrets, the adversary has the exact same view in the cases $b=0$ and $b=1$.)
@@
It is sufficient to prove that $\mathcal{A}_{n}$’s inputs, namely $D$, $Q_{n}$, and $R_{n}$, follow the same joint conditional distribution, conditioned to the event $\bar{E}$, regardless of whether $b=0$ or $b=1$.
@@
$\underset{}{\mathsf{Pr}}\left[{D,Q_{n},R_{n}\,|\,\bar{E},b}\right]=\,$ $\underset{}{\mathsf{Pr}}\left[{q_{n},r_{n}\,|\,D,Q_{n-1},R_{n-1},\bar{E},b}\right]\cdot\underset{}{\mathsf{Pr}}\left[{D,Q_{n-1},R_{n-1}\,|\,\bar{E},b}\right]$
@@
$\underset{}{\mathsf{Pr}}\left[{q_{n}\,|\,D,Q_{n-1},R_{n-1},\bar{E},b}\right]\cdot\underset{}{\mathsf{Pr}}\left[{r_{n}\,|\,q_{n},D,Q_{n-1},R_{n-1},\bar{E},b}\right]$
The first factor is independent of $b$ because $q_{n}$ is the output of $\mathcal{A}_{n-1}$, which only takes $D,Q_{n-1},R_{n-1}$ as inputs, all of which are among the conditioning terms already. The second factor is independent of $b$ because, given that $\bar{E}$ holds, the oracles behave the same whether $b=0$ or $b=1$. This is because striking queries are the only oracle inputs that could trigger responses with unequal distributions (i.e., correlated to $D$) in the cases $b=0$ and $b=1$. But if, instead, we rule these queries out by conditioning on $\bar{E}$, then the oracles instantiated when $b=0$ are perfectly interchangeable with the ones instantiated when $b=1$.
@@
The blank spaces (empty DMB cells, unmapped slices, etc.) are filled with equally-distributed uniformly-random noise. The spaces occupied by volume $V_{\ell}$, when $b=0$, are filled with oracle responses to queries containing one of $V_{\ell}$’s secrets. These responses are fresh randomnesses, which follow the same distribution as the noise filling the same spaces when $b=1$.
@@
The last step is to show that the position maps of the decoy volumes follow the same distribution in the two cases $b=0$ and $b=1$. By the second constraint on the access pattern, defined in Section 2.3.2, we get that slice allocation is triggered for the same LSIs of the decoy volumes in both the two cases. Even though some more slice allocations are performed on $V_{\ell}$’s LSIs when $b=0$, this does not impact the resulting observable distribution on the PSIs assigned to decoy LSIs. This is because slice allocation always takes a PSI randomly among the free ones, therefore the order in which the LSIs are mapped can be permuted freely without impacting the distribution. Thus, even in the case $b=0$, we can equivalently imagine that the decoy LSIs get all mapped before $V_{\ell}$’s ones, yielding the same distribution as when $b=1$. This concludes the proof.∎

### v2

@@ 4.1 Design
The permutation of indexes changes at every device instantiation, so that the mapping is not static. The size of these in-memory supporting data structures (array and bitfield) is $O(\frac{S_{\mathsf{D}}}{S_{\mathsf{P}}}\log\frac{S_{\mathsf{D}}}{S_{\mathsf{P}}})$, as usual. The lazy allocation technique is also what allows us to overcommit the total physical storage space: we can have the sum of the sizes of the logical volumes exceed the total physical storage space, as long as the sum of “actually used” spaces does not. However, it might be useful to intentionally limit this overcommitment (for the opened volumes) to decrease the risk of I/O errors and improve user experience. This can be done using the metadata field in the VMB ciphertexts, as explained in Section 6.5.
Handling slice conflicts.
Unsafe user operations could lead to accidental corruption of one or more volumes, resulting in the same PSI being assigned to different LSIs on different slice maps (see, e.g., Section 4.2). This would create a persistent ambiguity on volume instantiation, which would hamper any attempts from the upper layer (for example, as explained in Section 6.4) to restore data content. To avoid this, Shufflecake resolves any slice allocation conflict by duplicating and reassigning corrupted slices for every affected volumes. To be more precise:
  • When reading an entry $\psi$ from $\mathsf{SliceMap}_{i}[\sigma]$ for an LSI $\sigma$, a collision for $\psi$ is detected by seeing that the current element being initialised $\mathsf{ofld}[\psi]$ is already set to $\mathsf{occupied}$. That means that the physical slice with index $\psi$ has been corrupted: formerly, it belonged to volume $V_{i}$ (and, possibly, that was also corrupted from a volume of even higher order), but it has been mistakenly assigned also to a volume $V_{j}$ for some unknown $j<i$.
  • When this happens, do not attempt to resolve the collision by modifying $\mathsf{SliceMap}_{j}$. First, because by the time of this detection we might not know $j$, and second, because it might break plausible deniability in the multi-snapshot setting: If the corruption happens because an adversary has forced the user to open and write data to $V_{j}$ (as explained in Section 4.2, then the user is left alone and tries to repair the slice collision by changing $\mathsf{SliceMap}_{j}$, later at the next inspection the adversary sees that the content of $V_{j}$ has not changed, and yet $\mathsf{SliceMap}_{j}$ has, which would not be plausibly explainable.
  • Instead, the solution is to modify $\mathsf{SliceMap}_{i}$ by allocating a new slice with $\mathtt{NewSlice}(V_{i},\sigma)$, which also sets $\mathsf{SliceMap}_{i}[\sigma]$.
This has the disadvantage that the corruption of a few blocks due to operations performed on $V_{j}$ makes the whole slice unreadable by $V_{i}$ (because the freshly assigned slice contains garbage). However, remediating volume corruption is not a primary goal of Shufflecake, as this can only happen as a consequence of unsafe volume operations (see Section 4.2). The important point here is that this mechanism avoids the problem of allocation ambiguity persisting on the slice maps, thereby giving some hope for upper layer tools to repair the corrupted volumes. Possible improvements on this mechanism are described in Section 6.4. There, in particular, it is argued that by carefully cloning the content of the corrupted slice to the newly assigned slice, one can hope to recover from corruption at the block granularity rather than the slice granularity.
Device instantiation.
We are now ready to explain in detail how the process of device instantiation works. We assume a subroutine $\mathtt{HandleCorruption}$: this is the procedure responsible for handling the case of ambiguous slice assignment derived from a volume corruption. As explained above, in this reference description the corruption is handled by simply reassigning a new slice to the topmost corrupted volume. The full instantiation procedure is depicted in Algorithm 4. Notice that volume information is recovered ‘backward’, i.e., VMBs are decrypted starting from $N-1$ down to 00. However, once all VMBs (and related slice maps) are decrypted, the volumes are $\mathtt{open}$ed in forward order, meaning that $\mathsf{ofld}$ is populated by reading content from $\mathsf{SliceMap}_{0}$ to $\mathsf{SliceMap}_{N-1}$.
Algorithm 4 $\mathtt{Instantiate}$ ($\mathsf{device}$, $\mathsf{password}$)
1:read $\mathsf{salt}$ from $\mathsf{device}$
2:$\mathsf{KEK}:=\mathtt{KDF}(\mathsf{password},\mathsf{salt})$
3:for $j=0,\ldots,N_{\mathsf{MAX}}-1$
4:     read $(\mathsf{IV},{c})$ from $\mathsf{device}.\mathsf{DMB}.\mathsf{cell}[j]$
5:     $\mathsf{topvmk}:=\mathtt{Decrypt}({c},\mathsf{IV},\mathsf{KEK})$
6:     if $\mathsf{topvmk}\neq\bot$ then exit for $\triangleright$ Password correct: top volume found
7:next $j$
8:if $j=N_{\mathsf{MAX}}$ return “ERROR: wrong password.”
9:$N:=j+1$
10:global arrays $\mathsf{VMK}$, $\mathsf{VEK}$, $\mathsf{metadata}$ of size $N$, global var $\mathsf{numslices}$
11:$\mathsf{VMK}_{N-1}:=\mathtt{topvmk}$
12:for $i=N-1,\ldots,0$
13:     read $(\mathsf{IV},{c})$ from $\mathsf{device}.\mathsf{VMB}_{i}$
14:     $(\mathsf{VEK}_{i},\mathsf{prevVMK},\mathsf{thisnumslices},\mathsf{metadata}_{i}):=\mathtt{Decrypt}({c},\mathsf{IV},\mathsf{VMK}_{i})$
15:     if $i\neq 0$ then $\mathsf{VMK}_{i-1}:=\mathsf{prevVMK}$
16:     if $i=N-1$ then $\triangleright$ Sanity checks
17:          if $\mathsf{thisnumslices}>\lceil\frac{S_{\mathsf{D}}-S_{\mathsf{H}}}{S_{\mathsf{P}}}\rceil$ then return “ERROR: oversize volume.”
18:          $\mathsf{numslices}:=\mathsf{thisnumslices}$
19:     end if
20:     if $\mathsf{numslices}\neq\mathsf{thisnumslices}$ then return “ERROR: size mismatch.”
21:     global array $\mathsf{SliceMap}_{i}$ of size $\mathsf{numslices}$
22:     read $\mathsf{IV}^{\prime}$ and ${c}^{\prime}$ of size $\mathsf{numslices}$ from blocks after $\mathsf{device}.\mathsf{VMB}_{i}$
23:     $\mathsf{SliceMap}_{i}:=\mathtt{Decrypt}({c}^{\prime},\mathsf{IV}^{\prime},\mathsf{VEK}_{i})$
24:next $i$
25:global arrays $\mathsf{prmslices}$, $\mathsf{ofld}$ of size $\mathsf{numslices}$, global var $\mathsf{octr}$
26:$\mathsf{octr}:=0$
27:for $i=0,\ldots,\mathsf{numslices}-1$
28:     $\mathsf{prmslices}[i]:=i$
29:     $\mathsf{ofld}[i]:=\mathsf{free}$
30:next $i$
31:for $i=1,\ldots,\mathsf{numslices}-1$triangleright Fisher-Yates shuffle
32:     $j\raisebox{-1.0pt}{$\,\xleftarrow{\raisebox{-1.0pt}{$\scriptscriptstyle\$$}}\,$}\set{0,\ldots,i}$
33:     $\mathtt{swap}(\mathsf{prmslices}[i],\mathsf{prmslices}[j])$
34:next $i$
35:for $i=0,\ldots,N-1$triangleright Scan across all volumes
36:     for $\sigma=0,\ldots,\mathsf{numslices}-1$triangleright Scan all slice map of $V_{i}$
37:          $\psi\leftarrow\mathsf{SliceMap}_{i}[\sigma]$
38:          if $\mathsf{ofld}[\psi]=\mathsf{occupied}$ then
39:               $\mathtt{HandleCorruption}(V_{i},\psi,\sigma)$triangleright Corruption found; repair
40:          else
41:               $\mathsf{ofld}[\psi]:=\mathsf{occupied}$
42:          end if
43:      next $\sigma$
44:      create virtual block device $V_{i}$ of size $S_{\mathsf{L}}\cdot\mathsf{numslices}$
45:next $i$
46:return
4.1.3 Volume actions.
@@ 4.2 Operational Model
A simple observation shows how a legitimate-looking usage mode of Shufflecake actually entails a high risk of data corruption. If we do not open all $N$ existing volumes, and instead only open the ones we plan to use, we do not load all the slice maps in RAM, which leads to an incorrect reconstruction of the complete device’s bitfield ofld of free physical slices. The physical slices belonging to the still-closed volumes will be counted as free, and will therefore possibly be allocated to the open volumes during data write, which would then overwrite their content. This can only be avoided with certainty by always opening every volume, regardless of which ones we are going to use: if the password to a volume is not provided, Shufflecake has no way of detecting its existence. It then follows from the overcommitment of the physical storage space that we risk re-using its physical slices for some other volume.
However, mitigation does need to be perfect. It could be possible to reduce the risk of corruption when not opening all volumes by using some form of error-correction on the unopened volumes (hence sacrificing some space), and then trying to recover the volume if corruption happens. We discuss this in Section 6.4.
@@
In normal operating conditions, when not confronted with the adversary, the recommended course of action for the user is to unlock all the volumes present on the device, in order to prevent data corruption as explained before. The reason behind the design choice of chaining the volumes into a linked list to help the user in that regard: this way, the user is able to open all volumes by just providing the password to $V_{N-1}$. In our implementation, this is actually the mandated semantic of the $\mathtt{open}$ operation: the user only provides the password of the last volume they want to open, the previous ones are $\mathtt{open}$ed automatically. Other implementations might opt instead to ignore the $\mathsf{VMK}_{i-1}$ field in the volume header, and give more flexibility to the user, if aware of the risks entailed.
@@
When questioned by the adversary and forced to reveal passwords, the user must obviously not surrender more than $N-1$ of them (otherwise there would be nothing left to protect). Although irrelevant for the cryptographic security of the scheme, we stress that, in order for the user’s lie to be credible, they must only reveal the decoy passwords under a certain amount of pressure, or after some time has passed. Notice that a responsible and safe use of Shufflecake puts on the user the burden of being able recall quickly and reliably these decoy passwords, even under distress. This might be hard to get right given that the recommended course of action for the user in daily use is to only open the most hidden volume. It is up to the user to define the maximum number of volumes $N$ that makes them comfortable in this task. Shufflecake implementations might include additional features to aid the user in this sense, for example a function to check the password of a decoy volume without actually opening it, or even a puzzle which, with some random probability, prompts the user to also insert the password for a decoy volume when opening a hidden one.
@@
As we previously discussed, one big operational difference between Shufflecake (but also other solutions like, e.g., HiVE) and TrueCrypt is that with Shufflecake “the adversary does not know when she can stop questioning you”, because there is no way to prove that a given password unlocks all the existing volumes on a given device (unless it’s the password unlocking the $(N-1)$-th volume). In TrueCrypt, instead, there is either just a regular volume, or a decoy and a hidden volume. This distinction might be important in those scenarios where a user wants to avoid the possibility of looking uncooperative to a certain adversary. In such scenarios, the user might want to have the choice of surrendering all the volumes, and a method to convince the adversary that no other volumes exist, possibly using an additional “full disclosure password” that we will call safeword.
One simple way to implement this, even in Shufflecake, is to actually create all $N_{\mathsf{MAX}}$ volumes when initializing a device. Clearly, remembering $N_{\mathsf{MAX}}$ passwords would be quite cumbersome for the user, so the solution is to only remember $N+1$ passwords instead: those for the $N$ volumes that are actually desired, and the extra one for the last $N_{\mathsf{MAX}}$-th volume, which is going to be the safeword. In fact, the user will never need to open more than $N$ volumes for regular use. As discussed in Section 4.2, this might harm the consistency of the other, unopened volumes, but this is not important: by using the safeword, the user would still be able to convince the adversary that all volumes have been revealed due to the linkage between them and the ciphertext authentication in the headers. Analogously, in solutions like TrueCrypt, a simple way to implement a safeword is to actually always create a hidden volume, even if only a standard volume is desired.
@@ 4.3 Security
  • The symmetric encryption scheme will be replaced by an ideal cipher $\mathcal{E}$, mapping $\lambda$-bit keys and IVs to truly-random permutations over $\set{0,1}^{\lambda}$.
@@
Let us consider Experiment 2.1 under the constraints explained in Section 2.3.3, and let $D$ be the “challenge disk snapshot” provided to the adversary $\mathcal{A}$ by the game (i.e., either $D_{0}$ or $D_{1}$ according to the secret bit $b$). For the given $\mathcal{A}$, we will consider $N$, the decoy passwords $P_{0},...,P_{N-2}$, and the access patterns $O_{0}$ and $O_{1}$ as public parameters of the game instance.
@@
Then, let us define $Q$ to be the ordered sequence of all queries $\set{q_{i}}_{i}$ made by $\mathcal{A}$ to the oracles after receiving the challenge $D$, and define $n:=|Q|$. Also define $Q_{i}$ to be the sequence of queries $(q_{0},\ldots,q_{i-1})$ up to the $i$-th query. Analogously, let us define $R$ to be the ordered sequence of all responses $\set{r_{i}}_{i}$ returned by the oracles; also define $R_{i}$ to be the sequence of responses $(r_{0},\ldots,r_{i-1})$ up to the $i$-th response. We consider the execution of $\mathcal{A}$ as the execution of a sequence of single-query stateful adversaries, where the state is just the ‘history’ of the previous queries: $\mathcal{A}_{0}\left(D\right)\to q_{0}$, $\mathcal{A}_{1}\left(D,Q_{0},R_{0}\right)\to q_{1}$, and so on until $\mathcal{A}_{n-2}\left(D,Q_{n-2},R_{n-2}\right)\to q_{n-1}$ and $\mathcal{A}_{n-1}\left(D,Q_{n-1},R_{n-1}\right)\to b^{\prime}$.
Let $\mathsf{KEK}_{i}$, $\mathsf{VMK}_{i}$ and $\mathsf{VEK}_{i}$ be, respectively, the key-encryption key, the VMB key, and the volume encryption key of volume $V_{i}$. Rigorously speaking, in the security game, the values $P_{N-1},\mathsf{KEK}_{N-1},\mathsf{VMK}_{N-1}$, and $\mathsf{VEK}_{N-1}$ are only sampled if $b=0$. Let us instead consider them to be sampled anyway, and left unused in the case $b=1$. We denote by $\mathcal{S}$ the tuple $(P_{N-1},\mathsf{KEK}_{N-1},\mathsf{VMK}_{N-1},\mathsf{VEK}_{N-1})\in\set{0,1}^{4\lambda}$. Let us define the event $E_{i}$ as the event that either of $P_{N-1}$, $\mathsf{KEK}_{N-1}$, $\mathsf{VMK}_{N-1}$, or $\mathsf{VEK}_{N-1}$ appear in query $q_{i}$ (we say that query $q_{i}$ strikes). Finally, let us define $E:=E_{0}\cup\ldots\cup E_{n-1}$ the event that at least one query strikes. We will first prove two lemmata.
@@
$\underset{}{\mathsf{Pr}}\left[{E}\right]=\underset{}{\mathsf{Pr}}\left[{E\,|\,b=0}\right]=\underset{}{\mathsf{Pr}}\left[{E\,|\,b=1}\right]=\mathsf{negl}(\lambda)$. (The adversary can only guess one of the secrets of $V_{N-1}$ with negligible probability.)
@@
This tuple gets sampled uniformly at random from $\set{0,1}^{4\lambda}$, so its distribution does not depend on the public values. Since the oracles implement ideal cryptographic primitives, it follows that their outputs are statistically independent from their inputs; this does not just hold marginally for single input-output pairs, but jointly: any tuple of inputs is statistically independent from the tuple of corresponding outputs (for the ideal cipher $\mathcal{E}$, this only holds for the key inputs). In particular, $\mathcal{S}$ is statistically independent from the whole tuple $(D,R)$. Since $\mathcal{A}_{i}$’s inputs, namely the tuple $(D,Q_{i-1},R_{i-1})$, are a (randomised) function of $(D,R)$, we deduce that its output, namely $q_{i}$, must also be independent of $\mathcal{S}$.
Therefore, since $(P_{N-1},\mathsf{KEK}_{N-1},\mathsf{VMK}_{N-1},\mathsf{VEK}_{N-1})$ are uniform, the probability that $q_{i}$ contains, say, $P_{N-1}$, is $2^{-\lambda}$. By the union bound, $\underset{}{\mathsf{Pr}}\left[{E_{i}}\right]\leq 4\cdot 2^{-\lambda}$. Using the union bound again, we get:
$\underset{}{\mathsf{Pr}}\left[{E}\right]\leq\sum_{i=0}^{n-1}\underset{}{\mathsf{Pr}}\left[{E_{i}}\right]\leq n\cdot 4\cdot 2^{-\lambda}$
@@
$\underset{}{\mathsf{Pr}}\left[{\mathcal{A}\left(D\right)\to 1|b=0\wedge\bar{E}}\right]=\underset{}{\mathsf{Pr}}\left[{\mathcal{A}\left(D\right)\to 1|b=1\wedge\bar{E}}\right]$. (Unless she can guess one of $V_{N-1}$’s secrets, the adversary has the exact same view in the cases $b=0$ and $b=1$.)
@@
It is sufficient to prove that $\mathcal{A}_{n-1}$’s inputs, namely $D$, $Q_{n-1}$, and $R_{n-1}$, follow the same joint conditional distribution, conditioned to the event $\bar{E}$, regardless of whether $b=0$ or $b=1$.
@@
$\underset{}{\mathsf{Pr}}\left[{D,Q_{n\mbox{-}1},R_{n\mbox{-}1}|\bar{E},b}\right]\!=\!\underset{}{\mathsf{Pr}}\left[{q_{n\mbox{-}1},r_{n\mbox{-}1}|D,Q_{n\mbox{-}2},R_{n\mbox{-}2},\bar{E},b}\right]\!\cdot\!\underset{}{\mathsf{Pr}}\left[{D,Q_{n\mbox{-}2},R_{n\mbox{-}2}|\bar{E},b}\right]$
@@
$\underset{}{\mathsf{Pr}}\left[{q_{n-1}|D,Q_{n-2},R_{n-2},\bar{E},b}\right]\cdot\underset{}{\mathsf{Pr}}\left[{r_{n-1}|q_{n-1},D,Q_{n-2},R_{n-2},\bar{E},b}\right]$
The first factor is independent of $b$ because $q_{n-1}$ is the output of $\mathcal{A}_{n-2}$, which only takes $D,Q_{n-2},R_{n-2}$ as inputs, all of which are among the conditioning terms already. The second factor is independent of $b$ because, given that $\bar{E}$ holds, the oracles behave the same whether $b=0$ or $b=1$. This is because striking queries are the only oracle inputs that could trigger responses with unequal distributions (i.e., correlated to $D$) in the cases $b=0$ and $b=1$. But if, instead, we rule these queries out by conditioning on $\bar{E}$, then the oracles instantiated when $b=0$ are perfectly interchangeable with the ones instantiated when $b=1$.
@@
The blank spaces (empty DMB cells, unmapped slices, etc.) are filled with equally-distributed uniformly-random noise. The spaces occupied by volume $V_{N-1}$, when $b=0$, are filled with oracle responses to queries containing one of $V_{N-1}$’s secrets. These responses are fresh randomnesses, which follow the same distribution as the noise filling the same spaces when $b=1$.
@@
The last step is to show that the slice maps of the decoy volumes follow the same distribution in the two cases $b=0$ and $b=1$. By the second constraint on the access pattern, defined in Section 2.3.2, we get that slice allocation is triggered for the same LSIs of the decoy volumes in both the two cases. Even though some more slice allocations are performed on $V_{N-1}$’s LSIs when $b=0$, this does not impact the resulting observable distribution on the PSIs assigned to decoy LSIs. This is because slice allocation always takes a PSI randomly among the free ones, therefore the order in which the LSIs are mapped can be permuted freely without impacting the distribution. Thus, even in the case $b=0$, we can equivalently imagine that the decoy LSIs get all mapped before $V_{N-1}$’s ones, yielding the same distribution as when $b=1$. This concludes the proof.∎